# An Efficient Scene Coordinate Encoding and Relocalization Method

### v1

@@ V-A Comparison with State-of-the-Art
The 7-Scenes dataset comprises RGB-D images captured in seven indoor scenes using a Kinect RGB-D camera with a resolution of $640\times 480$. We train $f_{H}$ using the RGB frames and their ground-truth poses. Our approach is evaluated against several SOTA SCR systems and FM-based methods. Key metrics, including mapping time, map size, relocalization speed, and recall rate, are evaluated. The recall rate is defined as the percentage of test frames with pose errors below (1cm, 1). Results are summarized in Table I. Our single-mode relocalization achieves remarkable efficiency, running at 90, which is 1.6$\times$ faster than ACE, while also improving its recall by 6.4%. The sequence mode further delivers SOTA accuracy with high efficiency. By comparison, FocusTune and GLACE, both variants of ACE, have notable drawbacks: FocusTune relies on a 3D model and achieves only a modest accuracy gain (+3.9%), while GLACE sacrifices mapping and relocalization efficiency. In contrast, our system achieves a substantial recall improvement (+11%), while preserving ACE’s fast mapping time, compact map size, and relocalization speed.
@@
Sequence DSAC* ACE FocusTune GLACE Ours
Apt1_kitchen 87.1 84.9 88.0 93.3 97.5
Apt1_living 91.3 89.5 85.2 92.9 95.9
Apt2_bed 89.3 83.6 88.5 88.5 90.2
Apt2_kitchen 83.0 84.3 90.0 93.0 90.9
Apt2_living 83.6 89.1 89.4 91.1 91.6
Apt2_luke 66.5 76.4 77.1 82.7 84.8
Office1_gates362 73.1 73.6 76.2 94.6 81.1
Office1_gates381 62.3 68.2 70.0 85.2 78.8
Office1_lounge 75.8 70.6 69.1 82.3 81.7
Office1_manolis 70.9 72.9 71.4 84.4 77.2
Office2_5a 62.6 60.2 54.1 68.6 73.8
Office2_5b 58.3 71.4 61.2 84.0 85.9
@@
Sequence Median Error (cm / ) Recall (%)
ACE Ours $\Downarrow$ $\Uparrow$ ACE Ours $\Downarrow$ $\Uparrow$
GreatCourt 43 / 0.2 38 / 0.2 -5 / 0 6.7 9.5 +2.8
KingsCollege 28 / 0.4 25 / 0.4 -3 / -0 11.7 12.8 +1.1
OldHospital 31 / 0.6 27 / 0.6 -4 / 0 10.4 9.9 -0.5
ShopFacade 5 / 0.3 5 / 0.3 -0 / -0 85.4 86.4 +1.0
StMarysChurch 18 / 0.6 17 / 0.6 -1 / 0 20.2 23.0 +2.8
@@
The 12-Scenes dataset, collected in 12 scenes using an iPad color camera, provides RGB sequences at a resolution of $1296\times 968$. Like previous methods, we resize the image to $642\times 480$ during relocalization. We evaluate our system alongside several SOTA SCR approaches, with results summarized in Table II. Compared to the baseline (ACE), our methods improve relocalization recall by 8.6% and boost relocalization speed by 7. Although GLACE achieves slightly higher accuracy on this dataset, it comes at a significant cost: requiring 2.5$\times$ the mapping time, consuming 3$\times$ the memory for map storage, and operating at only half the relocalization speed of our approach. In Fig. 6, we compare the trajectory errors and recall rates of different algorithms for relocalization in two scenes.
@@
This dataset comprises five scenes from the old town of Cambridge, captured using a Google LG Nexus 5 smartphone. The captured videos were downsampled into low-frame-rate image sequences, leading to large motion gaps between consecutive frames. These gaps make feature tracking via optical flow challenging, so we evaluate only our single-mode relocalization on this dataset. The results, summarized in Table III, compare our system with ACE to highlight the benefits of incorporating sequential information during mapping. Metrics include the median pose error and the recall, defined as the percentage of test frames below a (10cm, 10°) pose error. Our methods reduce pose error by 3cm and improve recall by 1.4%, demonstrating the effectiveness of leveraging sequential information.
@@
TABLE IV: Comparison with other sequential SCR methods. We report the median pose error (cm / °) and the running speed.
Sequence Cai et al. [36] Tang et al. [37] KFNet [23] Ours
Chess 2.0 / 0.8 2.0 / 0.7 1.8 / 0.7 0.5 / 0.1
Fire 2.0 / 1.0 2.0 / 0.8 2.3 / 0.9 0.7 / 0.3
Heads 4.0 / 2.7 1.0 / 0.8 1.4 / 0.8 0.4 / 0.2
Office 3.0 / 0.8 3.0 / 0.8 2.5 / 0.7 0.9 / 0.2
Pumpkin 4.0 / 1.1 4.0 / 1.1 3.7 / 1.0 0.9 / 0.2
Kitchen 4.0 / 1.1 3.0 / 1.1 3.8 / 1.2 0.7 / 0.2
Stairs 18.0 / 3.9 4.0 / 1.6 3.3 / 0.9 2.4 / 0.7
@@
FPS - 15 20 59
V-B Comparison with Sequential Methods
We further compare our method with other SCR approaches that leverage sequential information, using the 7-Scenes dataset for evaluation. The median pose errors and running speeds of each system are summarized in Table IV. Our system outperforms all others, achieving the best accuracy across all seven scenes with 67% lower translation error and 67% lower rotation error compared to the second-best method. This improvement likely stems from the fact that the other methods encode all pixels, including non-informative or misleading regions, which can dilute the network’s encoding capacity and hinder mapping accuracy. Additionally, [36] introduces an extra module for dense feature extraction, while [37] and KFNet [23] rely on dense matching, which compromises efficiency. In contrast, our method employs a lightweight head for detecting sparse features and efficiently tracks them using optical flow, striking a better balance between accuracy and computational cost.
7-Scenes 12-Scenes
Recall (%) $\Uparrow$ Recall (%) $\Uparrow$
w/o KDH, SM, SR 55.2 77.1
w/o SM, SR 60.3 +5.1 80.4 +3.3
w/o SR 61.6 +6.4 81.4 +4.3
Ours (Full) 66.2 +11.0 85.7 +8.6
@@ V-C Ablation Study
We conduct an ablation study on the 7-Scenes and 12-Scenes datasets to evaluate the contributions of the proposed keypoint detection head (KDH), sequence-based mapping (SM), and sequence-based relocalization (SR). By sequentially removing these three modules, we calculate the recall rate, defined as the percentage of test frames below a (1cm, 1°) pose error. The results, presented in Table V, clearly demonstrate the effectiveness of the proposed components. Since our system builds upon ACE, the configuration “w/o KDH, SM, SR” corresponds to the original ACE. On the 7-Scenes dataset, KDH, SM, and SR improve the recall rate by 5.1%, 1.3%, and 4.6%, respectively, achieving a total improvement of 11.0%. Similarly, on the 12-Scenes dataset, these methods lead to an 8.6% increase in recall rate, further validating their impact.
@@ VI Conclusion
In this work, we propose an efficient and accurate SCR system. Our approach introduces a unified scene-agnostic module for both scene encoding and salient keypoint detection, enabling the system to prioritize encoding informative regions. To strengthen implicit triangulation, we associate keypoints across multiple images to effectively leverage sequential information. We further design two relocalization modes: the single-mode, which performs one-shot relocalization with high efficiency, and the sequence-mode, which incorporates temporal information to boost relocalization performance. Experimental results demonstrate that our methods significantly improve system performance, achieving remarkable accuracy and efficiency. Nevertheless, like other SCR systems, we use a CNN as the backbone, which limits scene point predictions to rely solely on information from a single patch. In future work, we plan to explore other architectures like ViT [38] to enhance SCR systems.

### v2

@@ V-A Comparison with State-of-the-Art
The 7-Scenes dataset captures $640\times 480$ resolution image sequences in seven indoor scenes. We train $f_{H}$ using the RGB frames and their ground-truth poses. Our approach is evaluated against several SOTA SCR systems and FM-based methods. Key metrics, including mapping time, map size, relocalization speed, and recall rate, are evaluated. The recall rate is defined as the percentage of test frames with pose errors below (1cm, 1). Results are summarized in Table I. Our single-mode relocalization achieves remarkable efficiency, running at 90, which is 1.6$\times$ faster than ACE, while also improving the recall by 6.4%. The sequence mode further delivers SOTA accuracy with high efficiency. By comparison, FocusTune and GLACE, both variants of ACE, have notable drawbacks: FocusTune relies on a 3D model and achieves only a modest accuracy gain (+3.9%), while GLACE sacrifices mapping and relocalization efficiency. In contrast, our system achieves a substantial recall improvement (+11%), while preserving ACE’s fast mapping time, compact map size, and relocalization speed.
@@
Sequence DSAC*[1] ACE[8] Focus- GLACE[10] Ours
Tune[9] Single Seq.
Apt1_kitchen 87.1 84.9 88.0 93.3 95.2 97.5
Apt1_living 91.3 89.5 85.2 92.9 90.9 95.9
Apt2_bed 89.3 83.6 88.5 88.5 88.9 90.2
Apt2_kitchen 83.0 84.3 90.0 93.0 84.3 90.9
Apt2_living 83.6 89.1 89.4 91.1 92.5 91.6
Apt2_luke 66.5 76.4 77.1 82.7 82.1 84.8
Office1_gates362 73.1 73.6 76.2 94.6 78.0 81.1
Office1_gates381 62.3 68.2 70.0 85.2 74.0 78.8
Office1_lounge 75.8 70.6 69.1 82.3 72.2 81.7
Office1_manolis 70.9 72.9 71.4 84.4 72.7 77.2
Office2_5a 62.6 60.2 54.1 68.6 68.6 73.8
Office2_5b 58.3 71.4 61.2 84.0 77.5 85.9
@@
Sequence Median Error (cm / ) Ensemble ($\times$4)
HLoc[3] GLACE[10] ACE[8] Ours ACE[8] Ours
GreatCourt 16/0.1 19/0.1 43/0.2 38/0.2 28/0.1 27/0.1
KingsCollege 12/0.2 19/0.3 28/0.4 25/0.4 18/0.3 17/0.3
OldHospital 15/0.3 17/0.4 31/0.6 27/0.6 25/0.5 24/0.5
ShopFacade 4/0.2 4/0.2 5/0.3 5/0.3 5/0.3 5/0.3
StMarysChurch 7/0.2 9/0.3 18/0.6 17/0.6 9/0.3 9/0.3
@@
  • In this table, “Ours" represents “Ours-Single."
@@
The 12-Scenes dataset provides $1296\times 968$ resolution image sequences across 12 scenes. Following prior works, we resize images to $642\times 480$ for relocalization. Table II presents the evaluation results against SOTA SCR methods. Compared to the baseline (ACE), our methods improve relocalization recall by 8.6% and boost relocalization speed by 7. Although GLACE achieves slightly higher accuracy on this dataset, it comes at a significant cost: requiring 2.5$\times$ the mapping time, consuming 3$\times$ the memory for map storage, and operating at only half the relocalization speed of our approach. In Fig. 6, we compare the trajectory errors and recall rates of different algorithms for relocalization in two scenes.
@@
This dataset comprises five scenes from the old town of Cambridge. The captured videos were downsampled into low-frame-rate image sequences, leading to large motion gaps between consecutive frames. These gaps make feature tracking via optical flow challenging, so we evaluate only our single-mode relocalization on this dataset. The results, summarized in Table III, show that Ours-Single reduces the average error of ACE by 3cm, demonstrating the effectiveness of leveraging sequential information in enhancing mapping accuracy. However, when using the ensemble method, the improvement of Ours-Single over ACE becomes minimal. This may be because the clustering process disrupts the original image sequence order, resulting in very limited sequence information available for Ours-Single.
@@
TABLE IV: Generalization validation.
Methods 7-Scenes 12-Scenes Cambridge
(%, $\uparrow$) (%, $\uparrow$) (cm / , $\downarrow$)
ACE 55.2 77.1 25 / 0.4
ACE + Ours 66.2 85.7 22 / 0.4
GLACE 61.6 86.7 14 / 0.3
GLACE + Ours 69.7 89.8 13 / 0.3
  • 1
    We report the percentage of test frames below a (1cm, 1) pose error on 7-Scenes and 12-Scenes, as well as the median pose errors on Cambridge Landmarks.
V-B Generalization Validation
Although our system builds upon ACE [8], the proposed methods are not dependent on the specific structure. To validate this, we also apply our methods to another SCR system, GLACE [10]. As shown in Table IV, our methods yield improvements on both ACE and GLACE, where "ACE + Ours" and "GLACE + Ours" denote the application of our methods to ACE and GLACE, respectively. The results indicate consistent performance gains across all three datasets. However, the improvement on Cambridge Landmarks is relatively smaller, likely due to its limited sequential information compared to the 7-Scenes and 12-Scenes datasets, as discussed in Section V-A. Overall, these experimental results highlight both the effectiveness and generalization capability of the proposed methods.
@@ V-B Generalization Validation
black Methods 7-Scenes 12-Scenes Recall (%) $\Uparrow$ Recall (%) $\Uparrow$ w/o KDH, SM, SR 55.2 77.1 w/o SM, SR 60.3 +5.1 80.4 +3.3 w/o SR 61.6 +6.4 81.4 +4.3 Ours (Full) 66.2 +11.0 85.7 +8.6
  • KDH, SM, and SR represent keypoint detection head, sequence-based mapping, and sequence-based relocalization, respectively.
@@ V-C Ablation Study
We perform an ablation study on the 7-Scenes and 12-Scenes datasets to assess the contributions of the proposed keypoint detection head (KDH), sequence-based mapping (SM), and sequence-based relocalization (SR). By sequentially removing these modules, we compute the recall rate, defined as the percentage of test frames with a pose error below (1cm, 1°). The results in Table V demonstrate the effectiveness of the proposed components. Since our system builds on ACE, the “w/o KDH, SM, SR” setting corresponds to ACE. On the 7-Scenes dataset, KDH, SM, and SR improve recall by 5.1%, 1.3%, and 4.6%, respectively, achieving a total improvement of 11.0%. On the 12-Scenes dataset, these methods yield an 8.6% gain, further validating their impact.
@@ VI Conclusion
In this work, we propose an efficient and accurate SCR system with a unified scene-agnostic module for scene encoding and keypoint detection, prioritizing informative regions. To enhance implicit triangulation, we associate keypoints across images to leverage sequential information. Additionally, we introduce two relocalization modes: a high-efficiency single-mode and a sequence-mode that integrates temporal cues for improved accuracy. Experimental results demonstrate that our methods significantly improve system performance, achieving remarkable accuracy and efficiency. Nevertheless, like other SCR systems, we use a CNN as the backbone, which limits scene point predictions to rely solely on information from a single local patch. In contrast, ViT [42] can capture long-range dependencies between image patches and leverage image global information to some extent. Therefore, for future work, we suggest training a scene-agnostic ViT encoder to extract feature vectors, enhancing the SCR system’s ability to leverage global image information.